OpenStudy (anonymous):

what is the distance between the z-intercept from the x-intercept in the equation ax+by+cz+d=0

OpenStudy (anonymous):

the z intercept will occur when x and y are both 0, so we get: cz+d = 0 z = -(d/c) the x intercept is when y and z = 0, so we get: ax +d = 0 x = -(d/a) using the pythagorean theorem we obtain: \[distance = \sqrt{(-\frac{d}{c})^2+(-\frac{d}{a})^2} = d\sqrt{\frac{1}{c^2}+\frac{1}{a^2}}\]

OpenStudy (anonymous):

tats it.... full ans is nt displayd...

OpenStudy (anonymous):

thts right answer riya

OpenStudy (anonymous):

k k thank u....